Circular No. 5938

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 S. N. Shore, Indiana University, South Bend; S. Starrfield and P.

Hauschildt, Arizona State University; R. González-Riestra, IUE Observatory, Vilspa; and G. Sonneborn, NASA/Goddard Space Flight Center, report: “Observations of Nova Cas 1993 with the IUE Satellite have confirmed the formation of dust (IAUC 5925, 5936) in the nova ejecta during the deep optical decline that is now underway (IAUC 5934). Spectra obtained on Feb. 18.1 and 19.5 UT show that the 120–200-nm portion of the spectrum has declined by more than a factor of ten since Feb. 11.2 UT with virtually no change in the absorption line opacity; the entire wavelength interval has been uniformly depressed. The longer-wavelength portion of the spectrum, 220–340 nm, exhibits depression by more than a factor of 20 in comparison to Feb. 11.5, with a broad peak in the extinction near Mg II 280 nm. The Feb. 18.3 long-wavelength spectra now show the C II 232-nm multiplet in emission, one of the few changes in the line spectrum during the past two weeks. There is no evidence for either enhanced absorption near 220 nm or for a steep short-wavelength rise. Therefore, the observed extinction, from the circumstellar dust that has just formed in the nova ejecta, does not resemble typical interstellar dust. The most likely explanation is that these grains are larger than normal interstellar dust. If the drop in the emitted radiation in the ultraviolet and optical is compensated by an increase in the infrared radiation as expected (cf. Gehrz 1988, Ann. Rev. Astron. Astrophys. 26, 377), the integrated infrared flux longward of 1 μm should reach of order 8 × 10^{−8} erg cm^{−2} s^{−1} (based on the ultraviolet flux depression). We have used E(B − V) = 0.5 to correct for the interstellar (not circumstellar) extinction. Before the decline began, the 120–340-nm continuum contained the bulk of the luminosity. The bolometric luminosity of the nova has been constant to within ∼ 10 percent from 1993 Dec. 15 to 1994 Feb. 11. It likely still is, except that it is now hidden by the enhanced extinction produced by the increasingly opaque circumstellar dust.”

PSR 0329+54 In the item by Dagkesamansky and Shitov on IAUC 5930, it should

have been made clear that the possibility of a 1110-day periodicity has been studied earlier (Demiansky and Proszynski 1979, Nature 282, 383; Cordes and Downs 1985, Ap. J. Suppl. 59, 343; Bailes, Lyne and Shemar 1993, ASP Conf. Ser. 36, 19).

1994 February 21 Brian G. Marsden